A stamp collection consists of 5, 13, and 18 cent stamps. The number of 13 cent stamps is one less than twice the number of 5 cent stamps. The number of 18 cent stamps is five more than the number of 13 cent stamps. The total value of all the stamps is $2.60. Find the number of 13 cent stamps in the collection
Answer by uma(370)

Let the number of 5 cent stamps be x
Then the number of 13 cent stamps = 2x - 1
The number of 18 cent stamps = 2x - 1 + 5
= 2x + 4
Value of 5 cent stamps = 5x cents
value of 13 cent stamps = 13(2x-1)
= 26x - 13 cents
value of 18 cent stamps = 18(2x + 4)
= 36x + 72 cents
Total value is 5x + 26x - 13 + 36x + 72 = 260
==> 67x + 59 = 260
==> 67x = 260 - 59
==> 67x = 201
==> x = 201/67
= 3
so x = 3
Thus the number of 13 cent stamps = 2x - 1
= 6 - 1
= 5
Good Luck!!!